OpenStudy (anonymous):

Let f(x)=(x−5)2 . Find the average rate of change of f(x) with respect to x from x= −3 to x= −3 + h if h≠0 . Simplify your answer as much as possible. Let f(x)=(x−5)2 . Find the average rate of change of f(x) with respect to x from x= −3 to x= −3 + h if h≠0 . Simplify your answer as much as possible. @Mathematics

OpenStudy (anonymous):

i get h or 1. im sure im doing it wrong though.

OpenStudy (anonymous):

i also got -16 + h / -67 which is also wrong

OpenStudy (anonymous):

You have to differentiate f(x)

OpenStudy (anonymous):

That doesnt help me much...

OpenStudy (anonymous):

I'm assuming you mean: f(x) = (x−5)^2?

OpenStudy (anonymous):

yes. i just copied and pasted.

OpenStudy (anonymous):

I assume you could find the average rate of change by finding the rate of change at f(-3) and f(-3+h) and then add those two together, and divide the result by 2 to get the average

OpenStudy (anonymous):

and the rate of change (sometimes called acceleration) is found by differentiation

OpenStudy (anonymous):

I know how to solve it. I must be doing something wrong though because Im getting the wrong answers.

OpenStudy (anonymous):

How do you do the differentiation?

OpenStudy (anonymous):

i got 64-16h+h^2-64 / -3+h-3

OpenStudy (anonymous):

then i canceled things out

OpenStudy (anonymous):

f(x) = (x-5)^2 = x^2 - 10x + 25 f'(x) = 2x - 10 f'(-3) = 2(-3) - 10 = -16 f'(-3+h) = 2(-3+h) - 10 = -6 + 2h - 10 (f'(-3) + f'(-3+h))/2 = (-16 - 6 + 2h - 10)/2 = (-32 + 2h)/2 = -16 + h

OpenStudy (anonymous):

I'm not entirely sure this is the right approach, there might be another way to calculate the average. Are you able to check if it's the correct answer?

OpenStudy (anonymous):

It's the right answer according to my online hw website. I'm trying to figure out how you got it though because I did not solve it like that.

OpenStudy (anonymous):

Hehe good. Where does it go wrong for you?

OpenStudy (anonymous):

i got -16+h/-6 earlier. that was the closest i got to the right answer.

OpenStudy (anonymous):

the -6 would have went away if it was -3 plus 3 not -3 minus 3

OpenStudy (anonymous):

so thats why i had the -6 too

OpenStudy (anonymous):

If you want to, you can show me the steps you go through. Because I can't really see how you get that result

OpenStudy (anonymous):

hold on a sec

OpenStudy (anonymous):

how did u get 2x-10?

OpenStudy (anonymous):

Just to be clear, do you know what differentiation is?

OpenStudy (anonymous):

ummm kind of

OpenStudy (anonymous):

I think you'll need a pretty good understanding of that, to solve these kinds of questions. Here's the wikipedia article http://en.wikipedia.org/wiki/Differentiation_(mathematics). It explains what it is better than I can.

OpenStudy (anonymous):

Basically it's a way to find the rate of change in regards to x, if that makes sense

OpenStudy (anonymous):

yeah thats what i thought... but i thought you find the change by subtracting old from new on the top and bottom of the fraction and then dividing the fraction and/or canceling things out.

OpenStudy (anonymous):

Yes that would work if it was a linear equation, because then the rate of change would be constant.

OpenStudy (anonymous):

ok thanks. i think i got it now. haha

OpenStudy (anonymous):

So, a = (y2-y2)/(x2-x1) is for linear equations only hehe

OpenStudy (anonymous):

ups (y2-y1)/(x2-x1)*

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

o wait u never answered me how u got 2x-10

OpenStudy (anonymous):

hehe all right

OpenStudy (anonymous):

f(x) = (x-5)^2 = x^2 - 10x + 25, so now we can differentiate f(x). The notation for which is f'(x) (notice the ' sign between the f and the (x) )

OpenStudy (anonymous):

(There are other notations that your teacher might prefer, like dx/dy instead of f'(x)

OpenStudy (anonymous):

so to differentiate x^2 - 10x + 25 we write f'(x) = (x^2 - 10x + 25)' =

OpenStudy (anonymous):

And here we use rules for differentiating

OpenStudy (anonymous):

a^x differentiated becomes x*a^(x-1)

OpenStudy (anonymous):

so x^2 becomes 2*x^(2-1)

OpenStudy (anonymous):

which is 2x^1 = 2x

OpenStudy (anonymous):

o wow. i dont recall doing this in class at all.

OpenStudy (anonymous):

hehe i hope i'm not trying to teach you things you don't need to understand yet

OpenStudy (anonymous):

what are you going through in class at the moment?

OpenStudy (anonymous):

We're on Functions part C

OpenStudy (anonymous):

hmm that doesn't tell me much

OpenStudy (anonymous):

We just started doing rate of change Friday

OpenStudy (anonymous):

haha judging by this link http://www.ies.co.jp/math/java/calc/heihen/heihen.html, i think i just made it worse for you

OpenStudy (anonymous):

none of your links are working

OpenStudy (anonymous):

oh sorry, there's a comma at the end of the link, there you go http://www.ies.co.jp/math/java/calc/heihen/heihen.html

OpenStudy (anonymous):

i think you were on the right track

OpenStudy (anonymous):

it's basically (f(-3+h)-f(-3))/h

OpenStudy (anonymous):

i didn't realize that since you're only interesting in finding the average rate of change you can consider it a linear equation

OpenStudy (anonymous):

forgive me for misleading you, you will probably have to learn the other way of doing it, later on though

OpenStudy (anonymous):

it's still a = (y2-y1)/(x2-x1) so f(-3+h) corresponds to y2, f(-3) corresponds to y1, and h corresponds to x2-x1